OpenStudy (nuccioreggie):

What is the value of a in the equation 4a + 3(a + 1) = 24? 2 3 7 21

OpenStudy (secret66):

4a + 3(a + 1) = 24 4a + 3a + 3 = 24 7a + 3 - 3 = 24 - 3 7a / 7 = 21 / 7 = 3 The answer is 3!

OpenStudy (nuccioreggie):

ok can you help me with 2 or 3 more please

OpenStudy (nuccioreggie):

An equation is shown below: 3(4x + 5) = 12 Which of the following correctly shows the beginning steps to solve this equation? Step 1: 12x + 15 = 12 Step 2: 12x = −3 Step 1: 12x + 5 = 12 Step 2: 12x = 7 Step 1: 7x + 5 = 12 Step 2: 7x = 7 Step 1: 7x + 8 = 12 Step 2: 7x = 4